OpenStudy (anonymous):

log(x) + log(x-15) = 2

OpenStudy (anonymous):

u want what?

OpenStudy (anonymous):

Solve logx+log(x−15)=2.

OpenStudy (anonymous):

I don't know the answer

OpenStudy (anonymous):

can u help me

OpenStudy (anonymous):

log(x(x-15))=2, using log rule

OpenStudy (anonymous):

Use the identities of logarithms to rewrite it as: log(x(x-15))=2 Then raise everything to the 10 (I'm assuming you mean log base 10). x(x-15)=10^2 x^2-14x=100 x^2-14x-100=0 Use quadratic formula.

OpenStudy (anonymous):

x(x-15)=10^2......log has base 10

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem :P

OpenStudy (anonymous):

u are very kind

OpenStudy (anonymous):

Use the identities of logarithms to rewrite it as: log(x(x-15))=2 Then raise everything to the 10 (I'm assuming you mean log base 10). x(x-15)=10^2 x^2-15x=100 x^2-15x-100=0 x^2-20x+5x-100=0 x(x-20)+5(x-20)=0 (x-20)(x+5)=0 So, x=20 or -5